OpenStudy (anonymous):

Evaluate the following integral:

OpenStudy (anonymous):

\[\int\limits_{\pi/3}^{\pi/2}\frac{ 1+\sin^2x }{ \sin^2x }dx\]

OpenStudy (experimentx):

split up the fraction ... evaluate it separately.

OpenStudy (anonymous):

That's when I get stuck... \[\int\limits_{\pi/3}^{\pi/2}\frac{ 1 }{ \sin^2x } + \int\limits_{\pi/3}^{\pi/2}\frac{ \sin^2x }{ \sin^2x }\]

OpenStudy (anonymous):

How do I find out the integral of 1/sin^2x?

OpenStudy (experimentx):

= csc^2(x) = 1 + cot^2(x) ..

OpenStudy (anonymous):

Isn't it just cot^2x for the integral of csc^2x? Why do you have 1 in there?

OpenStudy (sirm3d):

\[\int\frac{1}{\sin^2x}\mathrm dx=\int\csc^2 x\mathrm dx=-\cot x\]

OpenStudy (experimentx):

sorry .. my bad!!

OpenStudy (anonymous):

Great thanks people.